# arXiv:2603.00663v1[cs.RO] 28 Feb 2026

### v1

@@ 1 Introduction
The MT-VRP generalizes the Traveling Salesman Problem (TSP) and is therefore NP-hard (Hammar and Nilsson 1999; Helvig, Robins, and Zelikovsky 2003). Although previous work has investigated simplified versions of the MT-VRP without capacity constraints (Philip et al. 2025b; Stieber and Fügenschuh 2022), the capacitated version has not been considered and is the focus of this paper. Since state-of-the-art exact VRP solvers are based on the branch-and-price method (Costa, Contardo, and Desaulniers 2019), our objective is to develop a branch-and-price approach for MT-VRP as well.
@@
Figure 2: Example where existing VRP dominance checks fail for the MT-VRP. Since each target has one time window in this example, we represent a target-window simply using the index of the target. We treat the depot as a fictitious target 0. (a) $\tau_{\textnormal{a}}$} is an optimal (i.e. minimum-distance) trajectory visiting the sequence of targets $\Gamma=(0,3,1,2)$, and $\tau_{\textnormal{a}}$$^{\prime}$ is an optimal trajectory visiting the sequence $\Gamma^{\prime}=(0,1,3,2)$. Conventional VRP dominance checks would conclude that $\Gamma$ dominates $\Gamma^{\prime}$, since $\tau_{\textnormal{a}}$} and $\tau_{\textnormal{a}}$$^{\prime}$ have the same cost [^2] While dominance checks technically compare reduced cost, which adds bias terms to the cost, the bias terms for $\Gamma$ and $\Gamma^{\prime}$ are equal in Fig. 2 (a), and the same goes for $\bar{\Gamma}$ and $\bar{\Gamma}^{\prime}$ in Fig. 2 (b).. (b) $\bar{\tau}_{\text{a}}$ optimally visits the sequence $\bar{\Gamma}=(0,3,1,2,4,0)$ obtained by appending 4 and 0 to $\Gamma$, and $\bar{\tau}_{\text{a}}^{\prime}$ optimally visits the sequence $\bar{\Gamma}^{\prime}=(0,1,3,2,4,0)$ obtained by appending 4 and 0 to $\Gamma^{\prime}$. $\bar{\tau}_{\text{a}}$ travels more distance than $\bar{\tau}_{\text{a}}^{\prime}$, so $\bar{\Gamma}$ is worse than $\bar{\Gamma}^{\prime}$. Thus $\Gamma$ does not dominate $\Gamma^{\prime}$.
@@ 2 Related Work
Prior work (Philip et al. 2025b) finds optimal solutions for the Multi-Agent MT-TSP without capacity constraints using a mixed-integer conic program (MICP), assuming that targets move along piecewise-linear trajectories. We make the same piecewise-linear assumption in our work, and our baseline is based on (Philip et al. 2025b). Note that, using the terminology from the previous section, the MICP from (Philip et al. 2025b) is compact. Additionally, heuristic algorithms exist for the multi-agent MT-TSP, such as (Mathew, Smith, and Waslander 2015; Li et al. 2019, 2023), which only allow interception of each target at a finite number of sample points. However, these heuristics have no guarantees on solution quality.
@@ 3 Problem Setup
Each target $i$ has a demand $d_{i}\in\mathbb{R}_{\geq 0}$ that must be met by exactly one agent. It can be visited by any agent during one of its time windows in $\{[\underline{t}_{i,1},\overline{t}_{i,1}],\dots,[\underline{t}_{i,n_{\text{win}}(i)},\overline{t}_{i,n_{\text{win}}(i)}]\}$, where $\underline{t}_{i,j},\overline{t}_{i,j}]\subseteq\mathbb{R}_{\geq 0}$ is the time window $j^{th}$ of target $i$, and $n_{\text{win}}(i)$ is the number of time windows of target $i$. In addition, each target $i$ moves along a trajectory $\tau_{i}:\bigcup\limits_{j=1}^{n_{\text{win}}(i)}[\underline{t}_{i,j},\overline{t}_{i,j}]\rightarrow\mathbb{R}^{2}$ and has a constant velocity within a time window, but possibly different velocities in different time windows. We assume that no target at any time moves faster than $v_{\text{max}}$.
@@ 4 Integer Linear Program (ILP) for MT-VRP
$\mathcal{E}_{\text{tw}}$ contains an edge from target-window $\gamma_{i,j}$ to $\gamma_{i^{\prime},j^{\prime}}$ if and only if $i\neq i^{\prime}$. This edge records the latest feasible departure time from $\gamma_{i,j}$ to $\gamma_{i^{\prime},j^{\prime}}$, denoted by ${\textnormal{LFDT}}(\gamma_{i,j},\gamma_{i^{\prime},j^{\prime}})$. This value333(Philip et al. 2025a) describes how to compute LFDT. specifies the maximum $t\in[\underline{t}_{i,j},\overline{t}_{i,j}]$ for which there exists a speed-admissible trajectory that intercepts $\gamma_{i^{\prime},j^{\prime}}$ after intercepting $\gamma_{i,j}$ at time $t$.
@@
We now formulate an ILP where a solution to the MT-VRP is described by a set of tours, $\mathcal{F}_{\text{sol}}$. For a tour $\Gamma$, let $\alpha(i,\Gamma)=1$ if $\Gamma$ contains a target-window $\gamma_{i,j}$ of target $i$ and $\alpha(i,\Gamma)=0$ otherwise. Let the set of all possible tours for the agents be $\bar{\mathcal{F}}=\{\Gamma_{1},\Gamma_{2},\dots,\Gamma_{|\bar{\mathcal{F}}|}\}$. For each $\Gamma_{k}\in\bar{\mathcal{F}}$, define a binary variable $\theta_{k}$, which equals 1 if $\Gamma_{k}$ is chosen to be included in $\mathcal{F}_{\text{sol}}$ and 0 otherwise. Our ILP for the MT-VRP is as follows:
$\underset{\displaystyle\{\theta_{k}\}_{\Gamma_{k}\in\bar{\mathcal{F}}}}{\mathrm{min}}\quad\sum\limits_{\Gamma_{k}\in\bar{\mathcal{F}}}c^{*}(\Gamma_{k})\theta_{k}$ (1a)
@@
$\sum\limits_{\Gamma_{k}\in\bar{\mathcal{F}}}\alpha(i,\Gamma_{k})\theta_{k}\geq 1\;\;\forall i\in\{1,\dots,{n_{\textnormal{tar}}}\}$ (1c)
[…]
(1a) minimizes the sum of costs of the selected tours. (1b) ensures that we select up to $n_{\textnormal{agt}}$} tours. (1c) ensures that each target is visited by at least one tour. (1d) constrains $\theta_{k}$ to be binary. Note that the capacity constraints are implicit in the above ILP, since a tour must fully satisfy the demands of all targets it visits.
A convex relaxation of ILP (1) consists of (1a)-(1c), but replaces (1d) with the constraint $\theta_{k}\geq 0$; as in prior work (Feillet 2010), we remove the upper limit on $\theta_{k}$ in the relaxation. Given a set $\mathcal{B}\subseteq\mathcal{E}_{\text{tw}}$, we define ILP-$\mathcal{B}$ as ILP (1), with the constraint that $\theta_{k}=0$ for every $\Gamma_{k}$ traversing an edge in $\mathcal{B}$. Let LP-$\mathcal{B}$ be the convex relaxation of ILP-$\mathcal{B}$.
@@ 5 Branch-and-Price with Relaxed Continuity
We first provide an overview of our approach and then focus on the pricing problem. At the start of BPRC, we generate an initial feasible solution $\mathcal{F}_{\text{inc}}$ with cost $c_{\text{inc}}$, using the method from Section 5.3. We call $\mathcal{F}_{\text{inc}}$ the incumbent and continually update it with the best solution found so far. We also initialize a set of tours $\mathcal{F}$, setting it equal to $\mathcal{F}_{\text{inc}}$; we continually add to $\mathcal{F}$ throughout the algorithm.
BPRC then constructs a branch-and-bound tree, where a node is a set $\mathcal{B}\subset\mathcal{E}_{\text{tw}}$ of disallowed edges, designed to eliminate fractional solutions from the relaxation of ILP (1). We expand nodes from a stack in a depth-first fashion. When expanding a node $\mathcal{B}$, we solve LP-$\mathcal{B}$ (traditionally called the master problem in branch-and-price) using column generation. That is, when we initially solve LP-$\mathcal{B}$, we only optimize over $\theta_{k}$ variables for tours $\Gamma_{k}\in\mathcal{F}$, traversing no edge in $\mathcal{B}$, to keep the problem size tractable, assuming all other $\theta_{k}$ values are zero. This restriction of LP-$\mathcal{B}$ is known as the restricted master problem (RMP). After solving the RMP using an off-the-shelf LP solver, we solve a pricing problem to add tours to $\mathcal{F}$ that may improve the optimal cost of the RMP. Section 5.2 describes our method of solving the pricing problem. We alternate between the RMP and the pricing problem until the pricing problem adds no tours to $\mathcal{F}$, implying we have an optimal solution to LP-$\mathcal{B}$.
Whenever we find a new best integer solution during column generation, we update $\mathcal{F}_{\text{inc}}$ and $c_{\text{inc}}$. Additionally, the first time we solve the RMP for branch-and-bound node $\mathcal{B}$, the LP solver may return infeasible. This means that either no feasible MT-VRP solution can be assembled from the tours in $\mathcal{F}$, or all such solutions traverse an edge in $\mathcal{B}$. In this case, we attempt to use the method from Section 5.3 to produce a feasible MT-VRP solution $\mathcal{F}_{\text{new}}$ traversing no edge in $\mathcal{B}$, set $\mathcal{F}\leftarrow\mathcal{F}\cup\mathcal{F}_{\text{new}}$, then solve the RMP again. If we do not find any $\mathcal{F}_{\text{new}}$, LP-$\mathcal{B}$ is infeasible, and we discard the branch-and-bound node $\mathcal{B}$.
@@ 5.1 Pricing Problem
The pricing problem seeks a set of tours $\mathcal{F}_{\text{price}}=\{{}^{1}\Gamma,{}^{2}\Gamma,\dots,{}^{n_{\text{found}}}\Gamma\}$ such that the RMP on $\mathcal{F}\cup\mathcal{F}_{\text{price}}$ has a smaller optimal cost than the RMP on $\mathcal{F}$. $n_{\text{found}}$ is the number of tours found when solving the pricing problem and is not a user-specified parameter. We find $\mathcal{F}_{\text{price}}$ by leveraging the optimal dual solution to the RMP, denoted as $(\lambda_{0},\lambda_{1},\dots,\lambda_{{n_{\textnormal{tar}}}})$, where $\lambda_{0}\in\mathbb{R}_{\leq 0}$ is the dual variable corresponding to the constraint (1b), and $\lambda_{i}\in\mathbb{R}_{\geq 0}$ is the dual variable corresponding to the $i^{th}$ constraint in (1c). Following (Feillet 2010), a tour $\Gamma\in\mathcal{F}_{\text{price}}$ can reduce the optimal cost of the RMP on $\mathcal{F}\cup\mathcal{F}_{\text{price}}$ only if $\Gamma$ has a negative reduced cost, which is defined as follows:
@@
Thus the pricing problem seeks tours satisfying (2). [^5] A tour $\Gamma$ already in $\mathcal{F}$ must have ${c^{*}_{\textnormal{red}}}(\Gamma)\geq 0$, because otherwise $\lambda_{0},\lambda_{1},\dots,\lambda_{n_{\textnormal{tar}}}$ would be dual infeasible for the RMP (Feillet 2010). However, in practice, due to floating-point arithmetic, we may compute a slightly negative reduced cost for $\Gamma$. Thus, as in prior work (Kohl et al. 1999), we seek tours with reduced cost less than $\epsilon$ to avoid repeatedly adding tours to $\mathcal{F}$, where in our implementation, $\epsilon=10^{-4}$.
@@ 5.2 BPRC Labeling Algorithm
For a partial tour $\Gamma$ to dominate another partial tour $\Gamma^{\prime}$, two conditions must be met. First, any feasible extension of $\Gamma^{\prime}$ (i.e., a sequence of target-windows that can be appended to it) must also be a feasible extension of $\Gamma$. Feasibility here means not violating capacity constraints or revisiting targets. Second, the minimum reduced cost to execute the extension of $\Gamma$ must be no greater than that from $\Gamma^{\prime}$. Our primary contribution is a novel and computationally efficient method for verifying this second condition for the MT-VRP.
We first explain the core idea before providing formal definitions. Assume the feasibility condition is met and that both $\Gamma$ and $\Gamma^{\prime}$ end at the same target-window, $\gamma_{i,j}$. For an arbitrary $t^{\prime}\in[\underline{t}_{i,j},\overline{t}_{i,j}]$, let $\tau_{\textnormal{a}}$$^{\prime}$ be an optimal trajectory that executes $\Gamma^{\prime}$ and intercepts $\gamma_{i,j}$ at space-time point $(q^{\prime},t^{\prime})$, where $q^{\prime}=\tau_{i}(t^{\prime})$. Suppose there exists a trajectory $\tau_{\textnormal{a}}$} that executes $\Gamma$ and intercepts $\gamma_{i,j}$ at its starting point $(q,t)$. Let $\lambda=\sum\limits_{i=1}^{{n_{\textnormal{tar}}}}\alpha(i,{\tau_{\textnormal{a}}})\lambda_{i}+\lambda_{0}$, and $\lambda^{\prime}$ be defined similarly for $\Gamma^{\prime}$. Clearly, $\Gamma$ dominates $\Gamma^{\prime}$ if for all $t^{\prime}$, there exists a corresponding $\tau_{\textnormal{a}}$} satisfying
@@
We divide each target-window $\gamma_{i,j}$ into segments $\xi_{i,j,1},\dots,\xi_{i,j,n_{\text{seg}}(\gamma_{i,j})}$, where $\xi_{i,j,k}=(\gamma_{i,j},[\underline{t}_{i,j,k},\overline{t}_{i,j,k}])$ is the $k$th segment for $\gamma_{i,j}$, and $n_{\text{seg}}(\gamma_{i,j})$ is the number of segments for $\gamma_{i,j}$. The depot only gets one segment, i.e. $n_{\text{seg}}(\gamma_{0})=1$. We determine $n_{\text{seg}}(\gamma_{i,j})$ for $\gamma_{i,j}\neq\gamma_{0}$ via a user-specified parameter $n_{\text{seg,tar}}$, which specifies the number of segments per target. We allocate segments to target-windows so that every target-window gets at least one segment, and target-windows with longer time windows get more segments. Appendix C provides the segment allocation procedure.
@@
Figure 3: (a) and (b) show two partial tours $\Gamma$ and $\Gamma^{\prime}$ ending with $\gamma_{1,2}$, where we want to check if $\Gamma$ dominates $\Gamma^{\prime}$. (c) We check dominance by comparing an upper-bounding trajectory $\tau_{\text{a,ub}}$ for $\Gamma$ and a lower-bounding trajectory $\tau_{\text{a,lb}}^{\prime}$ for $\Gamma^{\prime}$, as described in Section 5.2. (d) We strengthen our dominance check by dividing $\gamma_{1,2}$ into segments and applying the check from (c) for each segment; an example is shown for the 5th segment of $\gamma_{1,2}$, denoted as $\xi_{1,2,5}$.
@@
  • $\vec{b}$ is a binary vector with length $n_{\textnormal{tar}}$}, where $\vec{b}[i^{\prime}]=1$ if and only if $\Gamma$ visits target $i^{\prime}$, $\sigma+d_{i}>d_{\text{max}}$ or $t>\max\limits_{j^{\prime}\in\{1,2,\dots,n_{\text{win}}(i^{\prime})\}}{\textnormal{LFDT}}(\gamma_{i,j},\gamma_{i^{\prime},j^{\prime}})$. The latter two conditions mean that target $i^{\prime}$ cannot be intercepted after intercepting $\gamma_{i,j}$ at time $t$, due to capacity or speed constraints [^6] Including both visitation and reachability status of targets in a label is a well-known strategy for strengthening dominance checks in VRP variants, introduced by (Feillet et al. 2004).
@@
  • $\vec{g}_{\text{lb}}$ is a vector of length $n_{\text{seg}}(\gamma_{i,j})$, where $\vec{g}_{\text{lb}}[k]$ is the cost of a lower-bounding trajectory for $\Gamma$ intercepting $\xi_{i,j,k}$
  • $\lambda=\sum\limits_{i=1}^{{n_{\textnormal{tar}}}}\alpha(i,\Gamma)+\lambda_{0}$.

### v3

# Optimal Solutions for the Moving Target Vehicle Routing Problem via Branch-and-Price with Relaxed Continuity

@@ 1 Introduction
The MT-VRP generalizes the Traveling Salesman Problem (TSP) and is therefore NP-hard (Hammar and Nilsson 1999; Helvig, Robins, and Zelikovsky 2003). Although previous work has investigated simplified versions of the MT-VRP without capacity constraints (Philip et al. 2025b; Stieber and Fügenschuh 2022), the capacitated version has not been considered and is the focus of this paper. Since state-of-the-art exact VRP solvers are based on the branch-and-price method (Costa, Contardo, and Desaulniers 2019), we present a branch-and-price approach for the MT-VRP.
@@
Figure 2: Example where existing VRP dominance checks fail for the MT-VRP. Since each target has one time window in this example, we represent a target-window using the index of the target. We treat the depot as a fictitious target 0. (a) $\tau_{\textnormal{a}}$} is an optimal (i.e. minimum-distance) trajectory visiting the sequence of targets $\Gamma=(0,3,1,2)$, and $\tau_{\textnormal{a}}$$^{\prime}$ is an optimal trajectory visiting the sequence $\Gamma^{\prime}=(0,1,3,2)$. [^3] Agents in the MT-VRP can change their speed, which is why $\tau_{\textnormal{a}}$} and $\tau_{\textnormal{a}}$$^{\prime}$ can travel the same distance, even though $\tau_{\textnormal{a}}$} meets target 2 later than $\tau_{\textnormal{a}}$$^{\prime}$: $\tau_{\textnormal{a}}$} can simply arrive at its interception position, then stop and wait for target 2. Conventional VRP dominance checks would conclude that $\Gamma$ dominates $\Gamma^{\prime}$, since $\tau_{\textnormal{a}}$} and $\tau_{\textnormal{a}}$$^{\prime}$ have the same cost. [^4] While dominance checks technically compare reduced cost, which adds bias terms to the cost, the bias terms for $\Gamma$ and $\Gamma^{\prime}$ are equal in Fig. 4 (a), and the same goes for $\bar{\Gamma}$ and $\bar{\Gamma}^{\prime}$ in Fig. 4 (b). (b) $\bar{\tau}_{\text{a}}$ optimally visits $\bar{\Gamma}=(0,3,1,2,4,0)$ obtained by appending 4 and 0 to $\Gamma$, and $\bar{\tau}_{\text{a}}^{\prime}$ optimally visits $\bar{\Gamma}^{\prime}=(0,1,3,2,4,0)$, obtained by appending 4 and 0 to $\Gamma^{\prime}$. $\bar{\tau}_{\text{a}}$ travels more distance than $\bar{\tau}_{\text{a}}^{\prime}$, so $\bar{\Gamma}$ is worse than $\bar{\Gamma}^{\prime}$. Thus $\Gamma$ does not dominate $\Gamma^{\prime}$.
@@ 2 Related Work
Prior work (Philip et al. 2025b) finds optimal solutions for the Multi-Agent MT-TSP without capacity constraints using a mixed-integer conic program (MICP), assuming that targets move along piecewise-linear trajectories. We make the same piecewise-linear assumption in our work. Our baseline is based on (Philip et al. 2025b). Note that, using the terminology from the previous section, the MICP from (Philip et al. 2025b) is compact.
VRP with Floating Targets:
The VRP with Floating Targets (VRPFT) (Gambella, Naoum-Sawaya, and Ghaddar 2018) is a problem similar to the MT-VRP, but where the movements of the targets must also be planned, and the targets do not have time windows. (Gambella, Naoum-Sawaya, and Ghaddar 2018) applies branch-and-price to the VRPFT, though the pricing problem is solved using an off-the-shelf mixed-integer program solver. In other branch-and-price methods, it is common to instead use a labeling algorithm for the pricing problem (Costa, Contardo, and Desaulniers 2019). In contrast to (Gambella, Naoum-Sawaya, and Ghaddar 2018), our work uses a labeling algorithm, introducing a new dominance criterion for moving targets.
@@ 3 Problem Setup
Each target $i$ has a demand $d_{i}\in\mathbb{R}_{\geq 0}$ that must be met by exactly one agent. It can be visited by any agent during one of its time windows in $\{[\underline{t}_{i,1},\overline{t}_{i,1}],\dots,[\underline{t}_{i,n_{\text{win}}(i)},\overline{t}_{i,n_{\text{win}}(i)}]\}$, where $\underline{t}_{i,j},\overline{t}_{i,j}]\subseteq\mathbb{R}_{\geq 0}$ is the $j$th time window of target $i$, and $n_{\text{win}}(i)$ is the number of time windows of target $i$. Each target $i$ moves along a trajectory $\tau_{i}:\bigcup\limits_{j=1}^{n_{\text{win}}(i)}[\underline{t}_{i,j},\overline{t}_{i,j}]\rightarrow\mathbb{R}^{2}$ and has a constant velocity within a time window, but possibly different velocities in different time windows. We assume no target at any time moves faster than $v_{\text{max}}$.
@@ 4 Integer Linear Program (ILP) for MT-VRP
$\mathcal{E}_{\text{tw}}$ contains an edge from target-window $\gamma_{i,j}$ to $\gamma_{i^{\prime},j^{\prime}}$ if and only if $i\neq i^{\prime}$. This edge records the latest feasible departure time from $\gamma_{i,j}$ to $\gamma_{i^{\prime},j^{\prime}}$, denoted by ${\textnormal{LFDT}}(\gamma_{i,j},\gamma_{i^{\prime},j^{\prime}})$. This value specifies the maximum $t\in[\underline{t}_{i,j},\overline{t}_{i,j}]$ for which there exists a speed-admissible trajectory that intercepts $\gamma_{i^{\prime},j^{\prime}}$ after intercepting $\gamma_{i,j}$ at time tt.555(Philip et al. 2025a) describes how to compute LFDT.
@@
We now formulate an ILP where a solution to the MT-VRP is described by a set of tours, $\mathcal{F}_{\text{sol}}$. For a tour $\Gamma$, let $\alpha(i,\Gamma)=1$ if $\Gamma$ contains a target-window $\gamma_{i,j}$ of target $i$ and $\alpha(i,\Gamma)=0$ otherwise. Let the set of all possible tours for the agents be $\mathcal{S}=\{\Gamma_{1},\Gamma_{2},\dots,\Gamma_{|\mathcal{S}|}\}$. For each $\Gamma_{k}\in\mathcal{S}$, define a binary variable $\theta_{k}$, which equals 1 if $\Gamma_{k}$ is chosen to be included in $\mathcal{F}_{\text{sol}}$ and 0 otherwise. Our ILP for the MT-VRP is as follows:
$\underset{\displaystyle\{\theta_{k}\}_{\Gamma_{k}\in\mathcal{S}}}{\mathrm{min}}\quad\sum\limits_{\Gamma_{k}\in\mathcal{S}}c^{*}(\Gamma_{k})\theta_{k}$ (1a)
@@
$\sum\limits_{\Gamma_{k}\in\mathcal{S}}\alpha(i,\Gamma_{k})\theta_{k}\geq 1\;\;\forall i\in\{1,\dots,{n_{\textnormal{tar}}}\}$ (1c)
$\theta_{k}\in\{0,1\}\;\;\forall k\in\{1,\dots,|\mathcal{S}|\}$ (1d)
(1a) minimizes the sum of costs of the selected tours. (1b) ensures that we select up to $n_{\textnormal{agt}}$} tours. (1c) ensures that each target is visited by at least one tour. (1d) constrains $\theta_{k}$ to be binary. Capacity constraints are implicit in ILP (1), since a tour must fully satisfy the demands of all targets it visits.
A convex relaxation of ILP (1) consists of (1a)-(1c), but replaces (1d) with the constraint $\theta_{k}\geq 0$; as in (Feillet 2010), we remove the upper limit on $\theta_{k}$ in the relaxation. An optimal solution $\theta^{*}=(\theta_{1}^{*},\theta_{2}^{*},\dots,\theta_{|\mathcal{S}|}^{*})$ to ILP (1) is feasible for this relaxation, but not necessarily optimal: some fractional solution may be optimal instead. Thus, the branch-and-bound in Section 5 iteratively breaks ILP (1) into subproblems, each disallowing some set of edges $\mathcal{B}$ to eliminate fractional solutions. The aim is to generate some subproblem where $\theta^{*}$ is optimal for the relaxation. In particular, for a set $\mathcal{B}\subseteq\mathcal{E}_{\text{tw}}$, we define the subproblem ILP-$\mathcal{B}$ as ILP (1), with the constraint that $\theta_{k}=0$ for each $\Gamma_{k}$ traversing some $e\in\mathcal{B}$. Let LP-$\mathcal{B}$ be the convex relaxation of ILP-$\mathcal{B}$.
@@ 5 Branch-and-Price with Relaxed Continuity
We first provide an overview of our approach and then focus on the pricing problem. At the start of BPRC, we generate an initial feasible solution $\mathcal{F}_{\text{inc}}$ with cost $c_{\text{inc}}$, using the method from Section 5.3. We call $\mathcal{F}_{\text{inc}}$ the incumbent and continually update it to be the best solution found so far. We also initialize a set of tours $\mathcal{F}$, setting it equal to $\mathcal{F}_{\text{inc}}$; we continually add tours to $\mathcal{F}$ throughout the algorithm.
BPRC then constructs a branch-and-bound tree, where a node is a set $\mathcal{B}\subseteq\mathcal{E}_{\text{tw}}$ of disallowed edges. We expand nodes from a stack in a depth-first fashion. When expanding a node $\mathcal{B}$, we solve LP-$\mathcal{B}$ (traditionally called the master problem in branch-and-price) using column generation. That is, when we initially solve LP-$\mathcal{B}$, we only optimize over $\theta_{k}$ variables for tours $\Gamma_{k}\in\mathcal{F}$, traversing no edge in $\mathcal{B}$, to keep the problem size tractable, assuming all other $\theta_{k}$ values are zero. This restriction of LP-$\mathcal{B}$ is known as the restricted master problem (RMP). After solving the RMP using an off-the-shelf LP solver, we solve a pricing problem to add tours to $\mathcal{F}$ that may improve the optimal cost of the RMP. Section 5.2 describes our method of solving the pricing problem. We alternate between the RMP and the pricing problem until the pricing problem adds no tours to $\mathcal{F}$, implying we have an optimal solution to LP-$\mathcal{B}$.
Whenever we find a new best integer solution during column generation, we update $\mathcal{F}_{\text{inc}}$ and $c_{\text{inc}}$. Additionally, the first time we solve the RMP for branch-and-bound node $\mathcal{B}$, the LP solver may return infeasible. This means that every feasible MT-VRP solution that can be assembled from the tours in $\mathcal{F}$ traverses some edge in $\mathcal{B}$. If so, we attempt to use the method from Section 5.3 to produce a feasible MT-VRP solution $\mathcal{F}_{\text{new}}$ traversing no edge in $\mathcal{B}$, set $\mathcal{F}\leftarrow\mathcal{F}\cup\mathcal{F}_{\text{new}}$, then solve the RMP again. If we do not find any $\mathcal{F}_{\text{new}}$, LP-$\mathcal{B}$ is infeasible, and we discard the branch-and-bound node $\mathcal{B}$.
@@ 5.1 Pricing Problem
The pricing problem seeks a set of tours $\mathcal{F}_{\text{price}}=\{{}^{1}\Gamma,{}^{2}\Gamma,\dots,{}^{n_{\text{found}}}\Gamma\}$ such that the RMP on $\mathcal{F}\cup\mathcal{F}_{\text{price}}$ has a smaller optimal cost than the RMP on $\mathcal{F}$. $n_{\text{found}}$ is the number of tours found when solving the pricing problem and is not a user-specified parameter. We find $\mathcal{F}_{\text{price}}$ by leveraging the optimal dual solution to the RMP, denoted as $(\lambda_{0},\lambda_{1},\dots,\lambda_{{n_{\textnormal{tar}}}})$, where $\lambda_{0}\in\mathbb{R}_{\leq 0}$ is the dual variable corresponding to the constraint (1b), and $\lambda_{i}\in\mathbb{R}_{\geq 0}$ is the dual variable corresponding to the $i^{th}$ constraint in (1c). Following (Feillet 2010), $\mathcal{F}_{\text{price}}$ can reduce the optimal cost of the RMP only if $\mathcal{F}_{\text{price}}$ contains a tour $\Gamma$ with negative reduced cost, which is defined as follows:
@@
Thus the pricing problem seeks tours with negative reduced cost. [^7] A tour $\Gamma$ already in $\mathcal{F}$ must have ${c^{*}_{\textnormal{red}}}(\Gamma)\geq 0$, because otherwise $\lambda_{0},\lambda_{1},\dots,\lambda_{n_{\textnormal{tar}}}$ would be dual infeasible for the RMP (Feillet 2010). However, in practice, due to floating-point arithmetic, we may compute a slightly negative ${c^{*}_{\textnormal{red}}}(\Gamma)$. Thus, as in prior work (Kohl et al. 1999), we seek tours with ${c^{*}_{\textnormal{red}}}(\Gamma)<-\epsilon$ to avoid repeatedly adding tours to $\mathcal{F}$; in our implementation, $\epsilon=10^{-4}$.
@@ 5.2 BPRC Labeling Algorithm
For a partial tour $\Gamma$ to dominate another partial tour $\Gamma^{\prime}$, two conditions must be met. First, any feasible extension of $\Gamma^{\prime}$ (i.e., a sequence of target-windows that can be appended to it) must also be a feasible extension of $\Gamma$. Feasibility here means not violating capacity constraints or revisiting targets. Second, the minimum reduced cost over extensions of $\Gamma$ must be no greater than that from $\Gamma^{\prime}$. Our primary contribution is a novel and computationally efficient method for verifying this second condition for the MT-VRP.
We first explain the core idea before providing formal definitions. Assume the feasibility condition is met and that both $\Gamma$ and $\Gamma^{\prime}$ end at the same target-window, $\gamma_{i,j}$. For an arbitrary $t^{\prime}\in[\underline{t}_{i,j},\overline{t}_{i,j}]$, let $\tau_{\textnormal{a}}$$^{\prime}$ be an optimal trajectory that executes $\Gamma^{\prime}$ and intercepts $\gamma_{i,j}$ at space-time point $(q^{\prime},t^{\prime})$, where $q^{\prime}=\tau_{i}(t^{\prime})$. Suppose there exists a trajectory $\tau_{\textnormal{a}}$} that executes $\Gamma$ and intercepts $\gamma_{i,j}$ at its starting point $(q,t)$. Let $\lambda=\sum\limits_{i=1}^{{n_{\textnormal{tar}}}}\alpha(i,\Gamma)\lambda_{i}+\lambda_{0}$, and $\lambda^{\prime}$ be defined similarly for $\Gamma^{\prime}$. Clearly, $\Gamma$ dominates $\Gamma^{\prime}$ if for all $t^{\prime}$, there exists a corresponding $\tau_{\textnormal{a}}$} satisfying
@@
We divide each target-window $\gamma_{i,j}$ into segments $\xi_{i,j,1},\dots,\xi_{i,j,n_{\text{seg}}(\gamma_{i,j})}$, where $\xi_{i,j,k}=(\gamma_{i,j},[\underline{t}_{i,j,k},\overline{t}_{i,j,k}])$ is the $k$th segment for $\gamma_{i,j}$, and $n_{\text{seg}}(\gamma_{i,j})$ is the number of segments for $\gamma_{i,j}$. The depot only gets one segment, i.e. $n_{\text{seg}}(\gamma_{0})=1$. We determine $n_{\text{seg}}(\gamma_{i,j})$ for $\gamma_{i,j}\neq\gamma_{0}$ via a user-specified parameter $n_{\text{seg,tar}}$, which is the number of segments per target. We allocate segments to target-windows so that each target-window gets at least one segment, and target-windows with longer time windows get more segments. Appendix C gives the allocation formula.
@@
Figure 3: (a) and (b) show two partial tours $\Gamma$ and $\Gamma^{\prime}$ ending with $\gamma_{1,2}$. We want to check if $\Gamma$ dominates $\Gamma^{\prime}$. (c) We check dominance by comparing an upper-bounding trajectory $\tau_{\text{a,ub}}$ for $\Gamma$ and a lower-bounding trajectory $\tau_{\text{a,lb}}^{\prime}$ for $\Gamma^{\prime}$, as described in Section 5.2. (d) We strengthen our dominance check by dividing $\gamma_{1,2}$ into segments and applying the check from (c) for each segment; an example is shown for the 5th segment of $\gamma_{1,2}$, denoted as $\xi_{1,2,5}$.
@@
  • $\vec{b}$ is a binary vector with length $n_{\textnormal{tar}}$}, where $\vec{b}[i^{\prime}]=1$ if and only if $\Gamma$ visits target $i^{\prime}$, $\sigma+d_{i^{\prime}}>d_{\text{max}}$, or $t>\max\limits_{j^{\prime}\in\{1,2,\dots,n_{\text{win}}(i^{\prime})\}}\hskip-2.84544pt{\textnormal{LFDT}}(\gamma_{i,j},\gamma_{i^{\prime},j^{\prime}})$. The last two conditions mean target $i^{\prime}$ cannot be intercepted after intercepting $\gamma_{i,j}$ at time $t$, due to capacity or speed constraints. [^8] Including both visitation and reachability status of targets in a label is a well-known strategy for strengthening dominance checks in VRP variants, introduced by (Feillet et al. 2004).
@@
  • $\vec{g}_{\text{lb}}$ is a vector of length $n_{\text{seg}}(\gamma_{i,j})$, where $\vec{g}_{\text{lb}}[k]$ is the cost of a lower-bounding trajectory for $\Gamma$ intercepting $\xi_{i,j,k}$.
  • $\lambda=\sum\limits_{i=1}^{{n_{\textnormal{tar}}}}\alpha(i,\Gamma)\lambda_{i}+\lambda_{0}$.